The length of a rectangle is 3 times longer than its width. If the area of the rectangle is 24 square inches, then what is the perimeter of the rectangle?

Answer by ewatrrr(24785)

 
Hi,
.
A = Lw
(3w)w = 24in^2 (Tossing out the negative solution for unit measure)
3w^2 = 24
w^2 = 8
w = 2√2in and L = 6√2in L+=+3w
P = 2L + 2w = 16√2in